OpenStudy (anonymous):

Write the equations in standard form for the lines satisfying the following conditions. Containing points (-1, 2) and (3, -5)

OpenStudy (jhannybean):

first find the slope m. m= (y2 -y1)/(x2 - x1) (x1, y1) = (-1, 2) (x2,y2) = (3,-5)

OpenStudy (jhannybean):

m = (-5-2)/ (3 - (-1) = -7 / 4 = -7/4 Next you want to use point slope form to get your equation in slope intercept form. y - y1 = m (x - x1) y - 2 = -7/4 ( x- (-1)) y - 2 = -7/4 (x + 1) y - 2 = -7/4x - 7/4 y = -7/4x - 7/4 + 2 y = -7/4x + 1/4 Standard form = Ax + By = C y + 7/4x = 1/4

OpenStudy (anonymous):

Thank you very much! You're super duper helpful